OpenStudy (anonymous):

Simplify (5/8)/(3/8)?

ganeshie8 (ganeshie8):

\[\frac{5/8}{3/8}\] multiply both numerator & denominator with 8 \[ = \frac{8*5/8}{8*3/8}\] \[ = \frac{\cancel{8}*5/\cancel{8}}{8*3/8}\] \[ = \frac{5}{\cancel{8}*3/\cancel{8}}\] \[= \frac{5}{3}\]

OpenStudy (anonymous):

Yeah I got it :) It was a silly question, but I have a bad case of summer brain!

ganeshie8 (ganeshie8):

lol.. its not silly question. its really very very difficult to visualize this particular division... when u get time, try to visualize below once : (five eighths of a "something" divided by "three eighths of some other thing") = ("five of somethings" divided by "three of some other things")